OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking another red marble? 4/45 1/15 1/9 8/75

OpenStudy (anonymous):

I'm thinking 1/15

OpenStudy (anonymous):

1/15

OpenStudy (anonymous):

why isn't it 1/9?

OpenStudy (anonymous):

5/15 * 5/15 = 1/3 * 1/3 = 1/9 ?? idk

OpenStudy (anonymous):

Well there are 15 balls in all that you are picking out of and not nine balls

OpenStudy (anonymous):

but don't you have to reduce it down

OpenStudy (anonymous):

Yea that got me too